OpenStudy (anonymous):

The sum of three numbers is 12. The first number is twice the sum of the second and third. The third number is 5 less than the first. Find the number. The key is: first is 8, second is 1, third is 3 This is complicated if you have to show your work and not just guess and check. Set up the equations first.

OpenStudy (anonymous):

There aren't really equations. You just have to use logic. Use that Logic, process of elimination.. I'm not sure what level of math you are in though

OpenStudy (anonymous):

this is algebra 2

OpenStudy (anonymous):

if you set up equation I know how to solve

OpenStudy (anonymous):

You could maybe use probability. Are you learning about Cnr and Pnr?

OpenStudy (anonymous):

yes I can do

OpenStudy (anonymous):

I can solve you don't have to solve , you just help me set up equation

OpenStudy (anonymous):

umm okay i'm not sure if this is right but

OpenStudy (anonymous):

figure out how many sums of 3 DIFFERENT (whole) numbers add up to 12. so 12*11*10

OpenStudy (anonymous):

so thats your total possibilities and now u can narrow it down

OpenStudy (anonymous):

given the key

OpenStudy (anonymous):

\[\{a+b+c=12,a=-10+2 a+2 b,c=a-5\} \]\[\{\{a\to 8,b\to 1,c\to 3\}\} \]

OpenStudy (anonymous):

thankyou haha

OpenStudy (anonymous):

robtobey how yu get number 10 and 5 ,

OpenStudy (anonymous):

the key for check answer not the given

OpenStudy (anonymous):

I'm working on a response.

OpenStudy (anonymous):

I still don't get ,how you get -10?

OpenStudy (anonymous):

{a + b + c = 12, a = 2 (b + c) , c = a - 5} Replace c in the first two equations with (a-5) {-5 + 2 a + b = 12, a = 2 (-5 + a + b)} Expand the second equation above by multiplying through by 2 {-5 + 2 a + b == 12, a == -10 + 2 a + 2 b} Solve the above two equations for a and b Take the value of a and plug it into c = a - 5 to obtain the value of c I hope I got this correct.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome.